T.C. = 14,500 + 200 Q – 4Q² + 2 Q³, Demand: Q = 1,200 – 0.5 P 1. At what level of Q (please round-off this Q to the nearest unit) and at what Price would this monopoly maximize profits? 2. Please use the marginal cost mark-up approach to calculate the Learner Index and interpret/explain your answer.
Answer #1

- 402 +20 a 14.500 + 200 ML = die/de NC = KX 200 - Bet 6c hi 1200 - 0.5 P. 30:5P 12 buc L P = (1260- (0) /0.5 P = 246-25 TR =Hence at 0= 19 (after round oft) muncpely muximizes profit. P: 24w - 2002 - 19 P = 2400-2 (19) P - 246 - 38 p = 2362 1 Monape